OpenStudy (anonymous):

logarithms:

OpenStudy (anonymous):

\[(\log_{7}8)(\log_{6}7)(\log_{5}6)(\log_{4}5)(\log_{3}4)(\log_{2}3) \]

OpenStudy (anonymous):

\[(\ln 8/\ln 7)(\ln 7/\ln 6)(\ln 6/\ln 5)(\ln 5/\ln 4)(\ln 4/\ln 3)(\ln 3/\ln2)\] =\[\ln 8/\ln 2\] =3

OpenStudy (anonymous):

oh duh!

OpenStudy (anonymous):

base of power rule!

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

k i see what you did. k thanks! *fanned you!

OpenStudy (anonymous):

You better! :P